OpenStudy (anonymous):

Find the derivative of f(x) = 1/2 (e^x + e^-x)

OpenStudy (anonymous):

haha, i just did this in class today, just a second

OpenStudy (anonymous):

do you want how to get it or just the answer?

OpenStudy (anonymous):

If you want how to get it I will toss it into the software I have so you know step by step, just let me know. Heres the answer though: f'(x)=(e^(-x)(e^(2x)-1))/(2)

OpenStudy (anonymous):

Could you show me how you got it.

OpenStudy (anonymous):

sure, I just looked over it and we got it the same way, where is this problem from by the way?

OpenStudy (anonymous):

f(x)=(1)/(2)*(e^(x)+e^(-x)) The derivative of f(x) is equal to f'(x)=(d)/(dx) (1)/(2)*(e^(x)+e^(-x)). f'(x)=(d)/(dx) (1)/(2)*(e^(x)+e^(-x)) Multiply the rational expressions to get ((e^(x)+e^(-x)))/(2). (e^(x)+e^(-x))/(2) Find the derivative of the expression. (d)/(dx) (e^(x)+e^(-x))/(2) The chain rule states that the derivative of a composite function (f o g)' is equal to (f' o g)*g'. To find the derivative of ((e^(x)+e^(-x)))/(2), find the derivatives of each portion of the function and use the chain rule formula. (d)/(du) (u)/(2)*(d)/(dx) e^(x)+e^(-x) To find the derivative of (u)/(2), multiply the base (u) by the exponent (1), then subtract 1 from the exponent (1-1=0). Since the exponent is now 0, u is eliminated from the term. (d)/(du) (u)/(2)=(1)/(2) The derivative of e^(x) is e^(x). This is true because (d)/(dx) a^(x)=a^(x)ln(a). In this case, the base is e and ln(e)=1, so the derivative is e^(x). (d)/(dx) e^(x)+e^(-x)=e^(x)+(d)/(dx) e^(-x) To find the derivative of -x, multiply the base (x) by the exponent (1), then subtract 1 from the exponent (1-1=0). Since the exponent is now 0, x is eliminated from the term. (d)/(dx) -x=-1 Using the chain rule, the derivative of e^(-x) is e^(-x)*-1. e^(-x)*-1 Multiply e^(-x) by -1 to get -e^(-x). -e^(-x) The derivative of e^(x)+e^(-x) is e^(x)-e^(-x). (d)/(dx) e^(x)+e^(-x)=e^(x)-e^(-x) Replace the variable u with e^(x)+e^(-x) in the expression. (d)/(du) (u)/(2)=(1)/(2) Replace the variable u with e^(x)+e^(-x) in the expression. e^(x)-e^(-x) Form the derivative by substituting the values for each portion into the chain rule formula. =(1)/(2)*(e^(x)-e^(-x)) Multiply the rational expressions to get ((e^(x)-e^(-x)))/(2). (d)/(dx) (e^(x)+e^(-x))/(2)=(e^(x)-e^(-x))/(2) Factor out the GCF of e^(-1x) from the expression e^(x). (d)/(dx) (e^(x)+e^(-x))/(2)=(e^(-1x)(e^(2x))-e^(-x))/(2) Factor out the GCF of e^(-1x) from the expression -e^(-x). (d)/(dx) (e^(x)+e^(-x))/(2)=(e^(-1x)(e^(2x))+e^(-1x)(-1))/(2) Factor out the GCF of e^(-1x) from e^(x)-e^(-x). (d)/(dx) (e^(x)+e^(-x))/(2)=(e^(-1x)(e^(2x)-1))/(2) Remove the -1 from -1x. (d)/(dx) (e^(x)+e^(-x))/(2)=(e^(-x)(e^(2x)-1))/(2) The derivative of the function is (e^(-x)(e^(2x)-1))/(2). f'(x)=(e^(-x)(e^(2x)-1))/(2)

OpenStudy (anonymous):

?

OpenStudy (anonymous):

he asked how to get it, Ihave a software program that generate step by step...with detailed explanations. Makes it easier than typing it all..lol

OpenStudy (anonymous):

why not \[\frac{1}{2}(e^x-e^{-x})\]

OpenStudy (anonymous):

if that is what he means satellitethen it is just the integrall of sinh(x)

OpenStudy (anonymous):

see. I thought about that, but my instructore earlier today worked the same exact problem, and this is the direction he went too. not sure whats up.

OpenStudy (anonymous):

am i lost? maybe tired. but there is nothing easier than finding the derivative of \[e^x\] it is just \[e^x\] and likewise the derivative of \[e^{-x}\] is \[-e^{-x}\] so you are done right away

OpenStudy (anonymous):

this does ask for "derivative" right? not "integral"

OpenStudy (anonymous):

yes but e^x-e^-x/2 is an identity for sinh(x)

OpenStudy (anonymous):

even this: f(x) = 1/2 (e^x + e^-x), see I agree with you. satellite

OpenStudy (anonymous):

yes sir, derivative

OpenStudy (anonymous):

so you could do dy/dx if the 1/2 is out in front

OpenStudy (anonymous):

that is what i did, and then applied the chain rule

OpenStudy (anonymous):

yes but if you knew that the identity e^x-e^-x /2 = sinh(x).. you could just do the dy/dx(sinh(x)) = cosh(x)

OpenStudy (anonymous):

yes if you wanted to prove that it was true, this would be the proof

OpenStudy (anonymous):

stick it to the teacher... make her thinkyou know more than what you should lmfao

OpenStudy (anonymous):

s quick quesiton, is the way I did it right or wrong?

OpenStudy (anonymous):

it seems like one step to me i suppose you use the "chain rule" to find the derivative of \[e^{-x}\]

OpenStudy (anonymous):

i don't know it if is right or wrong but it is awfully long to read. the derivative of e^x is e^x the derivative of e^(-x) is -e^(-x) and the one half is a red herring

OpenStudy (anonymous):

hey sat, if you don't mind, tell me exactly how come its one step? by the way, we haven't even made it to the chain rule in derrivvatives this semester, I just used it anyway.

OpenStudy (anonymous):

nm, u just told mehaha

OpenStudy (anonymous):

is this waht you have? \[\frac {dy}{dx}[\frac{1}{2(e^x-e^-x)}\]

OpenStudy (anonymous):

if so you would have to use chain rule because you pull 1/2 out and then you are left with (e^x-e^-x)^-1

OpenStudy (anonymous):

yes, thats it I think, I done put the notes away. i just posted all that to help him

OpenStudy (anonymous):

we had another one like it, hold on ill pull it up, it was intense, but I don't think it was necessary to do everything we ddid in that one either.

OpenStudy (anonymous):

or was it\[\frac{dy}{dx}[\frac{1}{2}e^x-e^-x]\]

jimthompson5910 (jim_thompson5910):

you did it right lawnphysics since (e^(-x)(e^(2x)-1))/(2)=(e^x-e^(-x))/2 is an identity (verified by wolfram alpha) But outkast3r09 has a good point \[\Large \frac{1}{2}(e^x+e^{-x})=\cosh(x)\] and \[\Large \frac{d}{dx}(\cosh(x)) = \sinh(x)=\frac{1}{2}(e^x-e^{-x})\] So \[\Large \frac{d}{dx}\left(\frac{1}{2}(e^x+e^{-x})\right)=\frac{1}{2}(e^x-e^{-x})\]

jimthompson5910 (jim_thompson5910):

so if you can see this, then it saves a lot of work

OpenStudy (anonymous):

i see both, thank you, that just helps me more. came in here to answer a question andlearned something too lol

OpenStudy (anonymous):

^ yes and lawnphysics, this stuff isn't really taught until integrals but it would help to get reps in right now too instead of taking this beyond what it needs to be

OpenStudy (anonymous):

yeah, but my instructor is hard core he don't like us using stuff outside of what he has alreadt tuahgt

OpenStudy (anonymous):

and the derivative isn't that bad also pull out one half and simply use the additive property so you have 1/2(dy/dx[e^x)+dy/dx[e^-x) dy/dx[e^u]=e^u u' so your just going to get e^x-e^-x

OpenStudy (anonymous):

here ill show you how to get it

OpenStudy (anonymous):

i hate how slow this thing types, has anyone ever though about a chat service just to speed things uP?

OpenStudy (anonymous):

See what you think of this one outkast3r09

OpenStudy (anonymous):

here ill show you how to get it \[\frac{dy}{dx}[\frac{1}{2}(e^x-e^-x)=\frac{1}{2}[\frac{dy}{dx}e^x-\frac{dy}{dx}e^-x]\]

OpenStudy (anonymous):

\[\frac{dy}{dx}[e^u]=e^uu'\]

OpenStudy (anonymous):

then simply solve

OpenStudy (anonymous):

now i know why he did it the way he did, the directions specifically say to use product or quotient rule then rewrite algebraically and apply quotient rule

OpenStudy (anonymous):

but knowing easier ways makes me happy anyway haha

OpenStudy (anonymous):

@jimthompson i must be getting tired, but what is the work in finding the derivative of e^x?

OpenStudy (anonymous):

dy/dx[e^u] = e^u u'

jimthompson5910 (jim_thompson5910):

I was referring to the lengthy steps lawnphysics wrote out

OpenStudy (anonymous):

so e^x is simple e^x

OpenStudy (anonymous):

and e^-x will be -e^-x

OpenStudy (anonymous):

i didn't write them out, Calculus solved did

jimthompson5910 (jim_thompson5910):

oh i see

OpenStudy (anonymous):

oh ok. still don't even see the need for sinh or cosh. it is two seconds to solve. the only tiny sophistication is knowing that the derivative of \[e^{-x}\] is \[-e^{-x}\]

OpenStudy (anonymous):

because it's much easier to just know that it's sinh(x) because the dy/dx of it is simple... +cosh(x)

OpenStudy (anonymous):

i use it to quickly find steps if I ever get stuck on a problem, not to cheat like most people. it helps if your working at 3am and there is no one to call haha

OpenStudy (anonymous):

and in that case you know that it's going to equal simple 1/2(e^x+e^-x)

OpenStudy (anonymous):

hey sat, do you use any instant messenging services?

OpenStudy (anonymous):

@outcast, why is the derivative of sinh cosh? because the derivative you just found shows it

OpenStudy (anonymous):

no not really. google chat sometimes

OpenStudy (anonymous):

i just use trillian, i hate how slow this things types haha

OpenStudy (anonymous):

the derivatives of hyperbolic functions are positive while their inverses are not i believe

OpenStudy (anonymous):

btw for your question lawnphysics the easier way (to me) is to write your denominator as (t^2-1)^-1

OpenStudy (anonymous):

and use the product rule

jimthompson5910 (jim_thompson5910):

you mean chain rule?

OpenStudy (anonymous):

well, the thing you have to rememver outkast, is we have only learned basic der. in class right now we are stuck using the basics. i understand there are better ways to do it too. can't want to be allowed to. i think ders. are freaking crazy fun

OpenStudy (anonymous):

\[\frac{d}{dt}[\frac{xt-4}{t^2-x}]=\frac{d}{dt}[(xt-4)(t^2-x)^{-1}\]

OpenStudy (anonymous):

jim, you would have to use the chain ffor the second part

OpenStudy (anonymous):

i just personally despise the quotient rule

jimthompson5910 (jim_thompson5910):

i must be looking at the wrong part

jimthompson5910 (jim_thompson5910):

oh ok I see now

OpenStudy (anonymous):

here, i despised it so much, i just took pictures of the board because I didn't want to write it down...here is what we did...i will post them...haha

OpenStudy (anonymous):

does anyone want to see them?

OpenStudy (anonymous):

i think i'm alright so did you need help on that ?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

nope, i was just saying thats how I learned to do the other one the original poster solved. and thats why I went that way with it

OpenStudy (anonymous):

just having a discusion about it with ya :)

OpenStudy (jamesj):

Christ almighty. What a lot of heat for so little light. @Sat73 is the one with the neon light here.

OpenStudy (anonymous):

well i think that the question was different from the question we answered still

OpenStudy (anonymous):

thnx. i thought i was losing my mind

OpenStudy (anonymous):

people need to use the equation editor

OpenStudy (anonymous):

because if the e's are in the bottom the problem becomes much longer like his teachers